OpenStudy (anonymous):

solve the system by substitution method -x + 8y = 16 3x - 24y = 48

OpenStudy (anonymous):

I'd simplify the first equation for x = something. 8y-16 = x Then, in the second equation substitute the 8y-16 in where you see x.

OpenStudy (anonymous):

3(8y-16) - 24y = 48 solve this for y...then you can find x

OpenStudy (anonymous):

This problem has a unique issue... 24y - 48 -24y = 48 -48 = 48 This statement is false, so there is no solution. The lines are parallel.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You're welcome